OpenStudy (anonymous):

4x^2 – 9x – 8 with x = –3

OpenStudy (anonymous):

... how hard can it be to enter numbers in a calculator?

OpenStudy (anonymous):

I keep getting the wrong answer

OpenStudy (anonymous):

4(-3)^2 -9(-3) -8

OpenStudy (anonymous):

4(9) +27 -8

OpenStudy (anonymous):

4*(-3)^2 - 9*(-3) - 8. Since (-3)^2 = (-3)*(-3), then (-3)^2 = 9. Also, (-9)*(-3) = 27. (4*9) + 27 - 8

OpenStudy (anonymous):

55?

OpenStudy (anonymous):

ya its 55